OpenStudy (anonymous):

Solve for Y showing all steps -10+2=5y-1

OpenStudy (anonymous):

-2x+2=5y-1 add 1 to both sides -2x+3=5y divide y from both sides -2/5x+3/5=y

OpenStudy (anonymous):

Sorry I mean divide 5 from both sides*

OpenStudy (anonymous):

Okay, thank you. Now, what if x was 5, how would I solve for y?

OpenStudy (shadowlegendx):

Are you still confused with this question?

OpenStudy (anonymous):

Yes

OpenStudy (shadowlegendx):

\[-10+2=5y-1 \] Do like terms first -10 + 2 = ?

OpenStudy (anonymous):

If x was 5, you would solve for y by just plugging in 5 for x and doing the same steps I did before

OpenStudy (shadowlegendx):

I don't see an x in the original equation...are we doing a different one than -10+2=5y-1 ?